how to determine if (2,-1) are solutions to 4r+3s=5
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Sub the numbers for r and s, see if it equals 5.
Is r = 2 and s = -1, or vice versa? Try it both ways, since it's not defined.

Answer by stanbon(75887)

determine if (2,-1) are solutions to 4r+3s=5
---------------
Substitute and see if you get a true statement:
4*2 + 3(-1) = 5
8 - 3 = 5
5 = 5
-------
r = 2 and s = -1 is a solution pair.
===================================
Cheers,
Stan H.
-----------------------------------